# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ Abstract.
We prove that every irreducible component of the coarse Kollár-Shepherd-Barron and Alexeev (KSBA) moduli space of stable log Calabi–Yau surfaces is a finite quotient of a projective toric variety. This verifies a conjecture of Hacking–Keel–Yu. The proof combines tools from log smooth deformation theory, the minimal model program, punctured log Gromov–Witten theory and mirror symmetry.
@@ 1.1. Context and overview
The higher dimensional analogue of a stable curve, introduced by Kollár–Shepherd-Barron [68] and Alexeev [5], is a stable pair $(Y,B)$, consisting of a projective variety $Y$ and a $\mathbb{Q}$-divisor $B$ such that $(Y,B)$ has semi-log-canonical (slc) singularities, and satisfying the stability condition requiring $K_{Y}+B$ to be $\mathbb{Q}$-Cartier and ample – see [66] for a comprehensive treatment of the KSBA theory. The geometry of the KSBA moduli space of stable pairs has been investigated so far in a handful of cases, including toric and abelian varieties [6], K3 surfaces [7, 8, 10, 72], elliptic surfaces [20, 21], and for some surfaces of general type [9]. Most relevant to the present paper, Alexeev showed in [6] that for toric varieties the normalization of the coarse KSBA moduli space is toric, with associated fan the secondary fan defined as in [37].
@@
More generally, for maximal log Calabi–Yau varieties, it has been conjectured by Keel [63] around 2012 that the coarse KSBA moduli space shall still admit a finite cover by a toric variety. Furthermore, Hacking–Keel–Yu conjectured that such a toric variety should be defined by a generalized version of the “secondary fan” coming from mirror symmetry [50]. We prove this conjecture in dimension two, that is, for all polarized maximal log Calabi–Yau surfaces $(Y,D,L)$ consisting of a projective surface $Y$ over $\mathbb{C}$, with a reduced, anticanonical divisor $D\subset Y$, such that $(Y,D)$ has at worst log canonical singularities, and an ample line bundle $L$ on $Y$. Furthermore, $(Y,D)$ is required to be maximal, in the sense that $D$ is non-empty and singular.
@@
The KSBA moduli space for $(Y,D,L)$, denoted by $\mathcal{M}_{(Y,D,L)}$, is the closure in the moduli space of stable pairs of the locus of stable pairs deformation equivalent to $(Y,D+\epsilon\,C)$, where $C\in|L|$ and $10<\epsilon<\!\!<$ 1. It is a proper Deligne–Mumford stack, with a projective coarse moduli space $M_{(Y,D,L)}$. To investigate the geometry of $M_{(Y,D,L)}$ via mirror symmetry, we construct a “semistable mirror” to a maximal degeneration of $(Y,D,L)$. Such a degeneration is determined by a choice of a polyhedral decomposition $\mathscr{P}$ on a “Symington polytope” $P$ associated to $(Y,D,L)$. To obtain the semistable mirror, we first build a normal crossing surface with dual intersection complex $(P,\mathscr{P})$. The semistable mirror is then obtained as a formal smoothing of this normal crossing surface. We show that such a smoothing is an open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\to\Delta=\mathrm{Spf}\,\mathbb{C}\llbracket t\rrbracket$, given by a projective crepant resolution $\mathcal{X}_{\mathscr{P}}\to\overline{\mathcal{X}}_{\mathscr{P}}$ of a formal affine canonical 3-fold singularity $\overline{\mathcal{X}}_{\mathscr{P}}$. We then define the “secondary fan” $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$, as a coarsening of the Mori fan for the semistable mirror to $(Y,D,L)$. Our main result, Theorem 8.11, states:
@@
The coarse KSBA moduli space $M_{(Y,D,L)}$ for a polarized toric log Calabi–Yau surface $(Y,D,L)$ admits a finite cover
@@
where $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is the toric variety associated to the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$.
We provide an outline of the key results involved in the proof in what follows, after briefly explaining why it goes via mirror symmetry.
@@ 1.2. Motivation from mirror symmetry
Given a stable log Calabi–Yau surface $(Y,D+\epsilon\,C)$ with an ample line bundle $L$, let $U:=Y\setminus D$, and $s$ be the section of $L|_{U}$ defining $C|_{U}$. One can naturally associate to $(Y,D+\epsilon\,C)$ the non-compact Calabi–Yau 3-fold $Z$ given by the hypersurface with equation $xy=s$ in the total space of the rank two vector bundle $L|_{U}\oplus\mathcal{O}_{U}$ over $U$, where $x$ and $y$ are linear coordinates along the fibers. Said differently, $Z$ is the affine conic bundle over the non-compact Calabi–Yau surface $U$ branched over the curve $C|_{U}$. Heuristically, the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ can be viewed as a compactification of a “moduli space of complex structures” on $Z$. Mirror symmetry predicts an identification of $\mathcal{M}_{(Y,D,L)}$ with the “stringy extended Kähler moduli space” of a mirror Calabi–Yau 3-fold $\mathcal{X}$ to $Z$ [77]. More precisely, under mirror symmetry different maximal degenerations of $Z$ should correspond to different mirrors related by birational transformations. In this paper, working directly with $(Y,D,L)$ instead of $Z$, we provide a precise realization of this heuristic picture, where a mathematically rigorous definition of the “stringy extended Kähler moduli space of $\mathcal{X}_{\mathscr{P}}$” is given by the toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$.
@@
It is worth noting that in the situation when $Y$ is a toric variety, and $D$ is the toric boundary divisor, mirror symmetry is entirely combinatorial and it essentially reduces to a duality between fans and momentum polytopes. Let $\overline{\mathcal{X}}$ be the singular toric variety with associated fan $C(P)$, the cone over the momentum polytope $P$ of $(Y,D,L)$. Choosing a maximal regular subdivision $\mathscr{P}$ and refining $C(P)$ by adding cones over $\mathscr{P}$, one obtains a toric crepant resolution $\mathcal{X}_{\mathscr{P}}\to\overline{\mathcal{X}}$, the semistable mirror family to $(Y,D,L)$. By standard toric geometry, the Mori fan of this mirror family to $(Y,D,L)$ coincides with the secondary fan of $(Y,D,L)$ [25]. So, the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ for a polarized toric log Calabi–Yau surface can be understood combinatorially, either in terms of the secondary fan, or of the Mori fan of the mirror [6, 98]. The existence of a dense torus in $\mathcal{M}_{(Y,D,L)}$ follows in this case from the existence of the basis of $H^{0}(Y,L)$ given by the monomials $z^{m}$ on $(\mathbb{C}^{\star})^{2}=Y\setminus D$ indexed by the integral points $m\in P_{\mathbb{Z}}:=P\cap\mathbb{Z}^{2}$. The key feature of this basis is that, for a curve $C$ defined by an equation $\sum_{m\in P_{\mathbb{Z}}}a_{m}z^{m}=0$ with $a_{m}\in\mathbb{C}^{\star}$ for all $m\in P_{\mathbb{Z}}$, the pair $(Y,D+\epsilon\,C)$ is KSBA-stable. Indeed, since for any vertex $m^{\prime}$ of $P$, having $a_{m^{\prime}}\in\mathbb{C}^{\star}$ ensures that $C$ does not intersect the singular point of $D$ corresponding to $m^{\prime}$, it follows that $(Y,D+\epsilon\,C)$ has only slc singularities if $a_{m}\in\mathbb{C}^{\star}$ for all $m\in P_{\mathbb{Z}}$ [6].
@@
Obtaining a toric description of the KSBA moduli space for general non-toric log Calabi–Yau surfaces, in particular finding a dense torus first inside the KSBA moduli space, again amounts to finding a basis for $H^{0}(Y,L)$, having the analogous key feature as the monomial basis in the toric case. However, by contrast to the toric case, purely combinatorial tools no more suffice to do so. Nonetheless, we show that such a basis is given by “theta functions”, defined by (a version of) the Gross–Siebert intrinsic mirror symmetry construction [48, 49], applied to the semistable mirror to $(Y,D,L)$. In restriction to a torus $(\mathbb{C}^{\star})^{2}\subset Y\setminus D$, the theta functions are particular Laurent polynomials with positive integer coefficients. However, no elementary combinatorial description of these Laurent polynomials is known. To describe them, following [48, 49], we use enumerative geometry, more precisely, particular counts of rational curves in the semistable mirror, defined using the theory of punctured Gromov–Witten invariants of Abramovich–Chen–Gross–Siebert [2].
@@ 1.3. Outline of the paper
In §2, we investigate maximal degenerations of polarized log Calabi–Yau surfaces combinatorially, using real affine geometry. To do so, we consider a “nice” toric model of a polarized log Calabi–Yau $(Y,D,L)$, as in previous work of Engel–Friedman [31], which allows us to systematically describe an associated “Symington polytope” $P$ to $(Y,D,L)$ given by an integral affine manifold with singularities [91]. We define the notion of a good polyhedral decomposition on $P$, and we prove the existence of such decompositions in Theorem 2.14. A good polyhedral decomposition $\mathscr{P}$ on $P$ defines in particular a maximal degeneration of $(Y,D,L)$, whose central fiber has “intersection complex” $(P,\mathscr{P})$, as explained in §2.4.
To construct a “mirror” to such a maximal degeneration of $(Y,D,L)$, we first construct a normal crossing surface $\mathcal{X}_{\mathscr{P},0}$ with “dual intersection complex” $(P,\mathscr{P})$ in §4.2. The irreducible components of $\mathcal{X}_{\mathscr{P},0}$ are obtained by $\mathbb{Q}$-Gorenstein smoothings of toric surfaces with quotient Wahl singularities, using Lemmas 4.1 and 4.2 which show that there are no local to global obstructions for such smoothings. Finally, we appropriately glue these irreducible components to obtain the normal crossing surface $\mathcal{X}_{\mathscr{P},0}$, which has features similar to the central fiber of a Type III Kulikov degeneration in the context of K3 surfaces [35]. A key difference in our context is that the dual intersection complex of $\mathcal{X}_{\mathscr{P},0}$ is a disk, rather than a sphere. So, in §3 we introduce the general notion of an open Kulikov surface, as a normal crossing surface $\mathcal{X}_{0}$ with dual intersection complex a disk, and satisfying some additional properties – see Definition 3.7. After investigating properties of such surfaces, as well as of their affinizations, we show in §4 that for an appropriate gluing the surface $\mathcal{X}_{\mathscr{P},0}$ associated is a $d$-semistable open Kulikov surface. Our key results in Theorems 3.26 and 3.29 prove the following result for a general open Kulikov surface $\mathcal{X}_{0}$:
@@
Let $\mathcal{X}_{0}$ be a quasi-projective open Kulikov surface. Then, there exists a quasi-projective formal smoothing $\mathcal{X}\to\Delta$ of $\mathcal{X}_{0}$ such that the total space is smooth and $\mathcal{X}_{0}\subset\mathcal{X}$ is a reduced normal crossing divisor, that is $\mathcal{X}\to\Delta$ is a semistable degeneration. Moreover, $K_{\mathcal{X}}=0$ and the affinization morphism $\mathcal{X}\to\overline{\mathcal{X}}$ is a projective crepant resolution.
@@
We refer to such a smoothing $\mathcal{X}\to\Delta$ as an open Kulikov degeneration. By the Minimal Model Program, adapted to the formal set up [75], we conclude in §3.6 that such a Kulikov degeneration is a relative Mori dream space. Hence, we obtain a complete toric fan, the “MoriFan”, in $\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})\otimes\mathbb{R}$. Then, following [50], we define the “secondary fan” $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$ as a “coarsening” of this Mori fan, by identifying neighboring cones connected by “M1 flops” – for details see §6.4, and denote by $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ the corresponding toric variety.
@@
For a polarized log Calabi–Yau surface $(Y,D,L)$ with a choice of $(P,\mathscr{P})$, we refer to the open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ obtained by smoothing $\mathcal{X}_{\mathscr{P},0}$ as the “semistable mirror” of $(Y,D,L)$. We then investigate the “double mirror”, obtained by applying a version of the intrinsic mirror construction of Gross–Siebert [48] to the semistable mirror $\mathcal{X}_{\mathscr{P}}\to\Delta$. More generally, we construct an intrinsic mirror for any quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$. One subtlety in our setup is that $\mathcal{X}\rightarrow\Delta$ is only quasi-projective, while the intrinsic mirror construction of [48] applies to projective degenerations. Therefore, we first prove in Theorem 5.2 that the moduli spaces of punctured log Gromov–Witten invariants of an open Kulikov surface $\mathcal{X}_{0}$, defining the structure constants in the intrinsic mirror construction, are proper. The proof crucially uses the fact that we have a projective crepant resolution $\mathcal{X}\to\overline{\mathcal{X}}$, where $\overline{\mathcal{X}}$ is affine. Hence, analogously as in [48], we define a commutative and associative graded mirror algebra $\mathcal{R}_{\mathcal{X}}$ – see Theorem 5.5. Thus, taking the Proj of this algebra, we obtain an intrinsic mirror family to $\mathcal{X}\to\Delta$ a priori defined over the formal scheme $\mathrm{Spf}\,\mathbb{C}\llbracket NE(\mathcal{X}/\overline{\mathcal{X}})\rrbracket$, as in [48], where $NE(\mathcal{X}/\overline{\mathcal{X}})$ stands for the monoid of integral points in the cone of relative effective curve classes of $\mathcal{X}\rightarrow\overline{\mathcal{X}}$.
@@
One of our key results, Theorem 6.16, shows that, for any quasi-projective open Kulikov degeneration $\mathcal{X}\to\Delta$, only finitely many curve classes of punctured log curves contribute to each structure constant of the intrinsic mirror algebra. This shows:

### v3

@@ Abstract.
We prove that every irreducible component of the coarse Kollár-Shepherd-Barron and Alexeev (KSBA) moduli space of stable log Calabi–Yau surfaces admits a finite cover by a projective toric variety. This verifies a conjecture of Hacking–Keel–Yu. The proof combines tools from log smooth deformation theory, the minimal model program, punctured log Gromov–Witten theory and mirror symmetry.
@@ 1.1. Context and overview
The higher dimensional analogue of a stable curve, introduced by Kollár–Shepherd-Barron [74] and Alexeev [4], is a stable pair $(Y,B)$, consisting of a projective variety $Y$ and a $\mathbb{Q}$-divisor $B$ such that $(Y,B)$ has semi-log-canonical singularities, and satisfying the stability condition requiring $K_{Y}+B$ to be $\mathbb{Q}$-Cartier and ample – see [72] for a comprehensive treatment of the KSBA theory. The geometry of the KSBA moduli space of stable pairs has been investigated so far in a handful of cases, including toric and abelian varieties [5], K3 surfaces [7, 8, 10, 78], elliptic surfaces [20, 21], and for some surfaces of general type [9]. Most relevant to the present paper, Alexeev showed in [5] that for toric varieties the normalization of the coarse KSBA moduli space is toric, with associated fan the secondary fan defined as in [40].
@@
More generally, for maximal log Calabi–Yau varieties, Keel [69] conjectured around 2012 that the coarse KSBA moduli space should still admit a finite cover by a toric variety. Furthermore, Hacking–Keel–Yu conjectured that such a toric variety should be determined by a generalized version of the “secondary fan”, derived from mirror symmetry [53, Conjecture 1.2]. In this paper, we prove this conjecture in dimension two, that is, for all polarized maximal log Calabi–Yau surfaces $(Y,D,L)$. These log Calabi–Yau surfaces consist of a projective surface $Y$ over $\mathbb{C}$, equipped with a reduced, anticanonical divisor $D\subset Y$, such that $(Y,D)$ has at worst log canonical singularities, and an ample line bundle $L$ on $Y$. Additionally, $(Y,D)$ must be maximal, meaning that $D$ is non-empty and singular.
The KSBA moduli space for $(Y,D,L)$, denoted by $\mathcal{M}_{(Y,D,L)}$, is the closure within the moduli space of stable pairs of the locus of stable pairs deformation equivalent to $(Y,D+\epsilon\,C)$, where $C\in|L|$ and $0<\epsilon<\!\!<1$. By [38, 75], $\mathcal{M}_{(Y,D,L)}$ is a proper Deligne–Mumford stack with a projective coarse moduli space $M_{(Y,D,L)}$. To investigate the geometry of $M_{(Y,D,L)}$, we follow the general strategy outlined by Hacking–Keel–Yu in [53, Conjecture 8.13], which proposes to use a conjectural double mirror construction. To realize this proposal, we first construct a “semistable mirror” to a maximal degeneration of $(Y,D,L)$. This degeneration is determined by choosing a polyhedral decomposition $\mathscr{P}$ on a “Symington polytope” $P$ associated with $(Y,D,L)$.
@@
The semistable mirror is obtained by first constructing a normal crossing surface with dual intersection complex $(P,\mathscr{P})$, followed by a formal smoothing of this surface. We show that such a smoothing is an open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\to\Delta=\mathrm{Spf}\,\mathbb{C}\llbracket t\rrbracket$, given by a projective crepant resolution $\mathcal{X}_{\mathscr{P}}\to\mathcal{X}^{\mathrm{can}}_{\mathscr{P}}$ of a formal affine canonical 3-fold singularity $\mathcal{X}^{\mathrm{can}}_{\mathscr{P}}$. Applying the general construction of the “secondary fan” in [53, §2] to the semistable mirror, we obtain a complete fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$ as a coarsening of the Mori fan of $\mathcal{X}_{\mathscr{P}}\to\mathcal{X}^{\mathrm{can}}_{\mathscr{P}}$. Our main result, Theorem 8.13, states:
@@
The coarse KSBA moduli space $M_{(Y,D,L)}$ for a polarized log Calabi–Yau surface $(Y,D,L)$ is the image of a finite surjective morphism
@@
where $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$ is the projective toric variety associated to the secondary fan $\mathrm{Sec}(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$.
We outline the key results involved in the proof in what follows, after briefly explaining why it goes through mirror symmetry, as predicted by [53, Conjecture 8.13].
@@ 1.2. Motivation from mirror symmetry
A stable log Calabi–Yau surface $(Y,D+\epsilon\,C)$ naturally defines a log Calabi–Yau 3-fold $(\overline{Z},\partial\overline{Z})$ as follows. The 3-fold $\overline{Z}$ is obtained by blowing up $C\times\{0\}$ in $Y\times\mathbb{P}^{1}$. The anticanonical divisor $\partial{\overline{Z}}$ is defined as the strict transform in $\overline{Z}$ of the anticanonical divisor $(Y\times\{0\})\cup(D\times\mathbb{P}^{1})\cup(Y\times\{\infty\})$ in $Y\times\mathbb{P}^{1}$. The complement $Z:=\overline{Z}\setminus\partial\overline{Z}$ is then a non-compact Calabi–Yau 3-fold. The projection $Y\times\mathbb{P}^{1}\rightarrow Y$ induces a map $Z\rightarrow U:=Y\setminus D$, which is an affine conic bundle on $U$ that degenerates over the divisor $C\cap U$. Heuristically, the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ can be viewed as a compactification of a “moduli space of complex structures” on $Z$. Mirror symmetry then predicts that $\mathcal{M}_{(Y,D,L)}$ should be identified with the “stringy extended Kähler moduli space” of a mirror Calabi–Yau 3-fold $\mathcal{X}$ to $Z$ [82]. As $Z$ is a maximal log Calabi–Yau 3-fold, the homology class of the fiber of the expected SYZ fibration on $Z$ is uniquely determined as the vanishing torus associated to a 0-stratum of the boundary. Consequently, distinct maximal degenerations of $Z$ should correspond to different mirrors related by birational transformations. In this paper, we work directly with $(Y,D,L)$ instead of $Z$. We provide a precise realization of this heuristic picture, where the “stringy extended Kähler moduli space of $\mathcal{X}_{\mathscr{P}}$” is rigorously defined as the toric variety $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$.
@@
In the situation where $Y$ is a toric variety and $D$ is the toric boundary divisor, mirror symmetry becomes entirely combinatorial, essentially reducing to a duality between fans and momentum polytopes. Let $\mathcal{X}^{\mathrm{can}}$ be the singular toric variety with associated fan $C(P)$, the cone over the momentum polytope $P$ of $(Y,D,L)$. A maximal regular subdivision $\mathscr{P}$ of $P$ defines a fan refining $C(P)$ by adding cones over $\mathscr{P}$, and so a toric crepant resolution $\mathcal{X}_{\mathscr{P}}\to\mathcal{X}^{\mathrm{can}}$, the semistable mirror family to $(Y,D,L)$. By standard toric geometry, the Mori fan of this mirror family to $(Y,D,L)$ coincides with the secondary fan of $(Y,D,L)$ [26]. Consequently, the KSBA moduli space $\mathcal{M}_{(Y,D,L)}$ for a polarized toric log Calabi–Yau surface can be understood combinatorially, either through the secondary fan or via the Mori fan of the mirror [5, 102].
The presence of a dense torus in $\mathcal{M}_{(Y,D,L)}$ follows in this case from the existence of the basis of $H^{0}(Y,L)$ given by the monomials $z^{m}$ on $(\mathbb{C}^{\star})^{2}=Y\setminus D$ indexed by the integral points $m\in P_{\mathbb{Z}}:=P\cap\mathbb{Z}^{2}$. The key feature of this basis is that, for a curve $C$ defined by an equation $\sum_{m\in P_{\mathbb{Z}}}a_{m}z^{m}=0$ with $a_{m}\in\mathbb{C}^{\star}$ for all $m\in P_{\mathbb{Z}}$, the pair $(Y,D+\epsilon\,C)$ is KSBA-stable. Indeed, for any vertex $m^{\prime}$ of $P$, ensuring that $a_{m^{\prime}}\in\mathbb{C}^{\star}$ guarantees that $C$ avoids the singular point of $D$ corresponding to $m^{\prime}$. As a result, $(Y,D+\epsilon\,C)$ has only semi-log-canonical singularities when $a_{m}\in\mathbb{C}^{\star}$ for all $m\in P_{\mathbb{Z}}$ [5].
To obtain a toric description of the KSBA moduli space for general non-toric log Calabi–Yau surfaces, particularly in identifying a dense torus within the KSBA moduli space, one must construct a basis for $H^{0}(Y,L)$ that exhibits the same key feature as the monomial basis in the toric case. However, unlike the toric setting, purely combinatorial methods are insufficient for this task. Nonetheless, mirror symmetry suggests that such a basis should consist of “theta functions” which are constructed using the enumerative geometry of curves in a mirror space [44, 47, 48, 51, 53, 70]. In this paper, we construct a basis of theta functions via a variant of the Gross–Siebert intrinsic mirror symmetry construction [51, 52], applied to the semistable mirror to $(Y,D,L)$.
@@
When restricted to a torus $(\mathbb{C}^{\star})^{2}\subset Y\setminus D$, the theta functions become specific Laurent polynomials with positive integer coefficients. However, no elementary combinatorial description of these Laurent polynomials is currently known. Instead, following [51, 52], we describe them through enumerative geometry, specifically via particular counts of rational curves in the semistable mirror, defined using the theory of punctured Gromov–Witten invariants of Abramovich–Chen–Gross–Siebert [2].
@@ 1.3. Outline of the paper
In §2, we investigate maximal degenerations of polarized log Calabi–Yau surfaces combinatorially, using real affine geometry. To do so, we consider a “nice” toric model of a polarized log Calabi–Yau $(Y,D,L)$, as in previous work of Engel–Friedman [33]. This allows us to systematically describe an associated “Symington polytope” $P$ to $(Y,D,L)$ given by an integral affine manifold with singularities [95]. We introduce the notion of a good polyhedral decomposition on $P$, and we prove the existence of such decompositions in Theorem 2.16. A good polyhedral decomposition $\mathscr{P}$ on $P$ naturally defines a maximal degeneration of $(Y,D,L)$, whose central fiber has “intersection complex” $(P,\mathscr{P})$, as explained in §2.4.
To construct a “mirror” to such a maximal degeneration of $(Y,D,L)$, we first construct a normal crossing surface $\mathcal{X}_{\mathscr{P},0}$ with “dual intersection complex” $(P,\mathscr{P})$ in §4.2. The irreducible components of $\mathcal{X}_{\mathscr{P},0}$ arise from $\mathbb{Q}$-Gorenstein smoothings of toric surfaces with quotient Wahl singularities. Using Lemmas 4.1 and 4.2, we show that these smoothings have no local to global obstructions. We then appropriately glue these irreducible components together to obtain the normal crossing surface $\mathcal{X}_{\mathscr{P},0}$, which exhibits features analogous to the central fiber of a Type III Kulikov degeneration in the context of K3 surfaces [37]. A key difference in our context is that the dual intersection complex of $\mathcal{X}_{\mathscr{P},0}$ is a disk rather than a sphere. In §3, we introduce the general notion of an open Kulikov surface, defined as a normal crossing surface $\mathcal{X}_{0}$ with a dual intersection complex that is a disk, subject to additional properties – see Definition 3.7. After investigating properties of such surfaces, along with their affinizations, we establish in §4 that, with appropriate gluing, the associated surface $\mathcal{X}_{\mathscr{P},0}$ is a $d$-semistable open Kulikov surface.
Our key results in Theorems 3.19 and 3.22 prove the following result for a general open Kulikov surface $\mathcal{X}_{0}$:
@@
Let $\mathcal{X}_{0}$ be a quasi-projective open Kulikov surface. Then, there exists a quasi-projective formal smoothing $\mathcal{X}\to\Delta$ of $\mathcal{X}_{0}$ such that the total space is smooth and $\mathcal{X}_{0}\subset\mathcal{X}$ is a reduced normal crossing divisor, that is $\mathcal{X}\to\Delta$ is a semistable degeneration. Moreover, $K_{\mathcal{X}}=0$ and the affinization morphism $\mathcal{X}\to\mathcal{X}^{\mathrm{can}}$ is a projective crepant resolution.
@@
We refer to such a smoothing $\mathcal{X}\to\Delta$ as an open Kulikov degeneration. By the Minimal Model Program, adapted to the formal setup [80], we conclude in §3.6 that such a Kulikov degeneration is a relative Mori dream space. This yields a complete toric fan, known as the “MoriFan”, in $\mathrm{Pic}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\otimes\mathbb{R}$. Applying the general construction of [53, §2], we obtain the “secondary fan” $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ as a “coarsening” of this Mori fan, by identifying neighboring cones connected by “M1 flops” – for details see §6.4. The associated toric variety is denoted by $S_{\mathcal{X}_{\mathscr{P}}}^{\mathrm{sec}}$.
@@
For a polarized log Calabi–Yau surface $(Y,D,L)$ with a choice of $(P,\mathscr{P})$, we refer to the open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ obtained by smoothing $\mathcal{X}_{\mathscr{P},0}$ as the “semistable mirror” of $(Y,D,L)$. Following the general strategy proposed by Hacking–Keel–Yu in [53, Conjecture 8.13], we then investigate the “double mirror” obtained by applying an intrinsic mirror construction [44, 47, 51, 70] to the semistable mirror. Specifically, we employ a variant of the intrinsic mirror construction of Gross–Siebert [51] to the semistable mirror $\mathcal{X}_{\mathscr{P}}\to\Delta$. More generally, we construct an intrinsic mirror for any quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$. A key subtlety in this setting is that $\mathcal{X}\rightarrow\Delta$ is only quasi-projective, whereas the intrinsic mirror construction of [51] is formulated for projective degenerations. To address this, we first establish in Theorem 5.2 that the moduli spaces of punctured log Gromov–Witten invariants of an open Kulikov surface $\mathcal{X}_{0}$, which define the structure constants in the intrinsic mirror construction, are proper. The proof critically relies on the existence of a projective crepant contraction $\mathcal{X}\to\mathcal{X}^{\mathrm{can}}$, where $\mathcal{X}^{\mathrm{can}}$ is affine. Consequently, following the approach of [51], we define a commutative and associative graded mirror algebra $\mathcal{R}_{\mathcal{X}}$ – see Theorem 5.6. Taking the Proj of this algebra, we obtain an intrinsic mirror family to $\mathcal{X}\to\Delta$, initially defined over the formal scheme $\mathrm{Spf}\,\mathbb{C}\llbracket NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\rrbracket$, as in [51]. Here, $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ stands for the monoid of integral points in the cone of relative effective curve classes of $\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$.
@@
Using techniques introduced in [44, 47, 70], we show in Theorem 6.24 that, for any quasi-projective open Kulikov degeneration $\mathcal{X}\to\Delta$, only finitely many curve classes of punctured log curves contribute to each structure constant of the intrinsic mirror algebra. This shows: